OpenStudy (anonymous):

how would i graph y=4/3x

OpenStudy (bahrom7893):

do u mean y = (4/3)x or y = 4/(3x) ?

OpenStudy (anonymous):

c'mon man. really??

OpenStudy (anonymous):

(4/3)x sorry

OpenStudy (bahrom7893):

Okay its a straight line with a slope of 4/3

OpenStudy (anonymous):

4/3 is a slope. So, Make a dot, and then go up 4 and over three. make another dot. when Y=0, what is x? thats your starting dot.